OpenStudy (anonymous):

Write this equation in slope-intercept form, and then find the slope and y-intercept. -y=3x

OpenStudy (anonymous):

slope-intercept form = y=mx+c

OpenStudy (anonymous):

Where m = slope c=y intercept

OpenStudy (anonymous):

y=-3x+0 Slope =-3 intercept =0

OpenStudy (anonymous):

-y=3x Multiply by -1 y = -3x + 0 y inetrcept = 0 slop = -3

OpenStudy (anonymous):

y=mx+b m is the slope, b is the y-intercept -y=3x → y=-3x + 0 slope, m, is -3. y-intercept, b, is 0 In order to find the x-intercept, you must plug in 0 for y 0=-3x → x=0/-3 → x=0

OpenStudy (anonymous):

So how exactly did you get to the slope and y-intercept?

OpenStudy (anonymous):

I'm confused..

OpenStudy (anonymous):

This form is called slope-intercept form because just by looking at the equation you can immediately tell what the slope and the y-intercept are m= slope b= y-intercept (another way to find the y-intercept is to plug in 0 for x)

OpenStudy (anonymous):

Okay...So how did you get -3?

OpenStudy (anonymous):

-y=3x You can divide both sides by - -y/-= y 3x/- = -3x y=-3x

OpenStudy (anonymous):

Oh...Thanks! :D